40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Engineering group Lab elog, Page 1 of 5  Not logged in ELOG logo
ID Date Author Type Category Subject
  252   Tue Jul 27 14:32:30 2021 Rich AbbottElectronicsHelicoflex Enclosure105kHz FC Detector

Luis, Chub, Calum, Jordan, Dean, Rich

Today we filled neon into the new enclosure (Dxxx) that's destined for use with the A+ filter cavity length and alignment detector, plus the new DCPD preamplifiers.  The goal is to do a neon accumulation test over in the 40m lab.

Here is some related information:

  1. The enclosure, lid, Helicoflex gasket, A286 10-32 SHCS bolts (used 20), S5 titanium washers (used 38*), S5 titanium nuts (used 20) were all cleaned and baked in the 40m bake facility by Jordan as an initial condition. * there were two holes where the washer would not fit under the head of the SHCS
  2. We had on hand:
    1. Fresh bottle of neon plus hoses and regulators (loosen the regulator knob for minimum pressure, tighten the knob for more pressure to load)
    2. Torque wrench
    3. Allan adapters for torque wrench to go into SHCS
    4. 3/8 inch spanner to hold nuts
    5. Fresh glove bag
    6. Vacuum pump plus hoses
    7. Kapton tape to seal hose ports
    8. Measuring tape to judge how much volume of neon we used
    9. iPad to take pictures as we went
    10. Big adjustable wrench and allen key to change out regulator on neon bottle
    11. Circuit board to install inside
    12. Ribbon cable to attach the circuit board to the inner wall of the enclosure (need strain relief design to be sure these connectors don't fall off)
  3. Here's what happened:
    1. Luis and Rich staged all the stuff in the clean room in the jitter lab
    2. We put the internal circuit board into the enclosure
    3. We plugged in the ribbon cable but had a bit of difficulty with putting the strain relief screws into the connectors (need a different setup here).  We ended up using some 4-40 screws in the connector on the wall of the enclosure, plus some zip ties to anchor down the connector to the PCB
    4. We didn't have any shorting plugs for the external connectors to avoid ESD damage, so we were careful, but no guarantees.
    5. We placed the gasket in its groove, then placed the lid on top of the enclosure.  Before tightening the bolts at all, we put a rolled up lint free cloth between the lid and the enclosure body so the neon had access to the inside of the enclosure while filling
    6. We put the allen adapters and 3/8 inch wrench into the glove bag
    7. We put the enclosure plus lid and gasket into a glove bag and rolled up as much of the bag as we could while still allowing room to get our hands into the access ports
    8. We added the vacuum hose and neon hose into the ports on the glove bag and taped them closed.
    9. We evacuated the bag with the vacuum pump, and then filled the bag with neon to a bag volume of approximately a 24-inch sphere.
    10. We removed the lint free cloth, then blew more neon into the enclosure to be sure it was filled well.  At this point, we were not thinking that neon was lighter than air.  Our first potential mistake.
    11. We snugged down the bolts and then removed the enclosure from the bag.  
    12. We then went off to lunch (our second mistake) and came back later to do the final torquing.
    13. When we returned, we realized that we now didn't know how much neon may have diffused from the interior of the enclosure, thus the experiment was uncertain.
    14. We proceeded anyway, and found torqued the bolts initially in a star pattern to 30 inch-pounds.
    15. After seven loops, (the second of which we abandoned the star pattern as it was too hard to keep our heads straight) we were at metal to metal on the flange surfaces, and could no longer rotate the bolts at 30 inch pounds.

We are chalking our inconclusive results up to experience, and starting again tomorrow with a fresh gasket.  We will be sure to account for the boyancy of neon in our fill method, and to rig a better way of flushing the interior of the enclosure with neon.

 

  251   Wed May 19 09:24:32 2021 StephenProgressRTS for COCKeyence Microscope integrated into RTS Setup

StephenA, 20 April 2021

Finished integrating the RTS Microscope Mount Assembly for the Keyence VH-Z250T lens (D2000085 WIP).

Feature description:

  • Lens assembly may be fully connected to microscope, then attached to the mount.
  • Easy on/off by passing lens assembly upward on shaft, then tightening thumb screw.
  • Locking shaft collars are added for security, once the lens is in the final position on the shaft and the thumb screw has been tightened
  • Position of lens assembly is repeatable in height and rotation.
    • Lens shaft mount (affixed by single thumb screw) registers in height and rotation against a flats cut into the vertical shaft. This provides registration of rotation about shaft axis.
    • Shaft is affixed to cantilever arm using shaft mounting blocks, including one with a set screw registering the rotation of the shaft with respect to the structure.
  • Lens shaft height and lens height may be adjusted independently to a large range of positions, using a series of thumbscrew flats and a continuous shaft registration flat.
  • Translation stage (used to focus microscope) has a hard stop in the direction toward the optic surface.

Overview of procedure for assembling the mount:

  1. Cantilevered aluminum extrusion arm is preassembled to interface plates on an optical table. Interface plates should be spaced appropriately for the destination breadboard - the spacing on a one inch grid can be confirmed on an arbitrary optical table, but the layout should be confirmed with measurements of the application.
  2. Shaft interface plates are preassembled to arm at a length along the arm that is appropriate for application.
  3. Translation stage may be preassembled to arm, or mounted in situ. Translation stage free plate is slid out of the way to access 1/4-20 clearance holes used to affix translation stage base plate to the aluminum adapter plate.
  4. Arm should be mounted to application at this stage.
  5. Shaft should be mounted to translation stage in situ, via the below process.
    1. A set screw through the base of one of the shaft mount blocks is driven into the shaft's continuous upper flat, registering the shaft's rotation. The shaft should be clamped fully on the base, with the set screw meeting the flat - needs some slight loosening of the upper clamp to attain the correct rotation.
    2. The shaft is lightly clamped into the second base, with roughly the correct alignment and spacing, which are
    3. The mount bases are placed on the translation stage, and adjustments are made with the bases gently loosened to align to the vertically-held shaft and interface to the tranlation stage threaded hole spacing.
      1. If loosening any mount block clamps, the shaft must be held to prevent a sudden drop!
    4. Overall shaft height is set by the position of the mount blocks along the shaft. Once the rotation-registering set screw is in place, shaft height may be adjusted by loosening the mount block clamps, with the mount block bases still on the translation stage.
  6. Confirm height of shaft provides clearance from optic (check for interference every time!)
  7. Add upper shaft clamp to provide a third clamping location and vertical stop.

Overview of procedure for installing the microscope:

  1. Bring the microscope lens assembly upward onto the shaft from underneath.
  2. Tighten the thumb screw into the correct flat, currently the second from the bottom.
    1. Helps to make sure the thumb screw is just outside of the ID of the mount when you start the installation, so that you can tell when the thumb screw has extended into the flat and past the OD of the shaft.
    2. I prefer to turn the thumb screw 2 turns, then gently lower the thumb screw into contact with the bottom surface of the flat. This provides a height registration and constraint for the lens assembly. Once I feel the lens assembly resting on the flat of the shaft, I then tighten the thumb screw into the vertical surface of the flat and lock in the position.
    3. Hold the lens assembly by the main body with one hand during this operaiton, to avoid dropping the lens
  3. Add the bottom shaft collar, typically stored loose above the permanent top shaft collar, to provide a redundant vertical height restraint.
  4. Use the translation stage to make any final height adjustments required to bring the optic into position under the microscope and avoid interference.
  5. Remove lens cap only when ready to bring the optic into position!

Images of mount in various states:

  1. Mount ready to host microscope lens assembly - IMG_8613
  2. Array of flats used to host the lens assembly's thumbscrew (we currently are mounted in the second from bottom flat) - IMG_8614
  3. Tightening of thumb screw, from similar POV to previous images (slightly higher zoom) - IMG_8615
  4. Lens assembly mounted and fully secured to mount - IMG_8617
  5. Overview image of lens assembly in mount; taken before fully secured, as lower shaft clamp is missing - IMG_8616
  6. Microscope hosted on small wire cart, in corner adjacent to RTS table - IMG_8467
Attachment 1: IMG_8613.JPG
IMG_8613.JPG
Attachment 2: IMG_8614.JPG
IMG_8614.JPG
Attachment 3: IMG_8615.JPG
IMG_8615.JPG
Attachment 4: IMG_8617.JPG
IMG_8617.JPG
Attachment 5: IMG_8616.JPG
IMG_8616.JPG
Attachment 6: IMG_8467.JPG
IMG_8467.JPG
  250   Thu Apr 8 17:01:52 2021 StephenHow ToModal TestingTroubleshooting low vibrometer signal / absence of ring down

Regina's problem statement:

 I attached examples of two measurements I got and wanted to know if these look reasonable. I took 6 measurements total, and I attached the first and last measurements. The graph on the last page is a picture of the weighting step for reference.

I didn't see significant ring down in all my measurements. Is this to be expected? I thought since the baffle is now rigidly mounted, the vibration should go almost to zero. I was also getting some low amplitude noise throughout the entire frequency domain that didn't show up in the first couple measurements I took. I tried to reset the vibrometer like you mentioned but they were still present. Is this a problem?

Stephen's reply:

1) Vibrometer may need a higher diffusivity surface to improve signal level.

Replying to your absence of ring down in your measurements - I agree that it looks like the vibrometer output is not behaving well, for one reason or another. In the freely suspended case, I was thinking this was due to large yaw and pitch motion causing high signal variation. Given that the symptom occurs when the baffle is fixed, I think the likeliest reason is the low signal, due to the low scatter and highly specular surface finish of this baffle (aka shiny). One way to troubleshoot would be to attach a compliant, diffusely reflecting material to the surface - think a small square of the adhesive-coated part of a Post-It note, for example - then tune the focus of the vibrometer and see if the vibrometer's signal level bar improves. If the signal level improves, take a hammer-excited measurement, and see if you see any ring down. If this behaves as you might expect, you could generate mode shape data with your excitation roving around the surface of the baffle while your response is fixed - just one baffle point would need your diffuse Post-It square (I might go with a central, or near central, location).

2) Test Article is not well understood, so try measuring something that has been characterized before.

If you try playing with the diffusivity and focus but the signal level doesn't improve, or if you don't see any ring down still, try pointing the vibrometerat the suspension cage and exciting the suspension with the hammer - that should give some real signal regardless of the precise setup, and if that gives similarly mystifying results, let me know and we can think a little harder about what might be going on. I feel pretty confident that between these two tests, you will find an answer.

3) Add another transducer (ie. a witness accelerometer) for comparison.

Another way to support your understanding of your setup (and a good practice) would be to mount the accelerometer to the suspension cage, adjacent to the baffle mounting brackets, or even to the mounting brackets themselves. This accelerometer would supply a witness to the low frequency resonances of the cage, which you may excite during your measurements, and might also provide some insights to the baffle panel resonances (rigid coupling with lower modal mass = smaller vibrations, but likely still above the noise floor of the accel) supporting your eventual successful vibrometer measurements.

4) Notes about mounting an accelerometer.

Mounting would involve collecting a ~[1mm x 1mm x 1mm] chunk of beeswax, spreading the beeswax onto the face of the accelerometer opposite the cable (I like to press it with the outer, flat-ish surface of my thumb nail to spread), and pressing the beeswaxed face onto a flat surface - think a 5 second push with all of your arm strength, which should create a thin layer with plenty of tackiness to hold the accel in any orientation. If it doesn't seem like it could hold for a day, then you might need more beeswax, or more force to create that thin layer. Note: the main way that an accel can go from useful to not is to experience a shock event, so I would recommend that you use some kapton tape to affix the accelerometer cable to the suspension cage - this will strain relieve your cabling and provide a fall restraint, and the potential for the cabling to influence the measurement is minimal here because it is remote from the baffle.

Attachment 1: regina_troubleshooting_GraphsBKHammer.pdf
regina_troubleshooting_GraphsBKHammer.pdf regina_troubleshooting_GraphsBKHammer.pdf regina_troubleshooting_GraphsBKHammer.pdf
  249   Tue Nov 24 12:31:05 2020 Marie K. SummaryBS BRDsLHO BRDs monitoring

LHO BRDs:

I didn't observe frequency drifts during the month of assembly and monitoring in the Optics Lab. This is not expected from our experiments in the Modal Lab, but it makes the preparation for the sites BRDs easier.

BOUNCE: In anticipation to the frequency drifts, I had tuned the BRDs on the lower side of the target frequency. But the tuning didn't drift so I changed the masses last week for the bounce mode in order to be in the 1% target. The jumps in the curves below are due to retuning (on November 5th for BRD1, on November 20th for the other ones).

Roll: I didn't retune the Roll modes after assembly on October 20th. In the last days, I was experimenting with different ways to excite them (see pictures attached), so this is probably the cause of the slight drifts that we see.

 


 

Attachment 3: Shaker_plate.jpg
Shaker_plate.jpg
Attachment 4: Shaker_electronics.jpg
Shaker_electronics.jpg
  248   Tue Nov 24 11:09:18 2020 Marie K.SummaryBS BRDsLHO BRDs ready for C&B

This morning I handed off the parts for the 5 LHO BRDs to Bob Cottingham at LLO for Clean&Bake (https://ics.ligo-la.caltech.edu/JIRA/browse/clean-10214)

The parts for each BRD are in separate containers (see pictures attached).

Here are the masses of each component in order to reassemble the BRDs after C&B:

  Roll   Bounce
  Copper Screw Washer Total [g] Copper Screw Washer Total [g]
BRD1 6.355 0.724 0.052 7.131 8.836 0.855 0.2 9.891
BRD2 6.393 0.839 0.05 7.282 8.816 0.844 0.2 9.86
BRD3 6.403 0.851 0.05 7.304 8.812 0.832 0.2 9.844
BRD4 6.372 0.85 0.05 7.272 8.88 8.49 0.2 17.57
BRD5 6.389 0.844 0.05 7.283 8.904 0.864 0.2 9.968
Attachment 1: 5_BRDS_C&B.jpg
5_BRDS_C&B.jpg
Attachment 2: Detail_1BRD_container.jpg
Detail_1BRD_container.jpg
Attachment 3: Load_C&B_1135.jpg
Load_C&B_1135.jpg
  247   Thu Nov 19 15:55:45 2020 StephenProgressVacuumNuts and Rod for A+ FC Tube Support Stand

Stephen A, 2020 Nov 19

Ordered and received PO 75-S492380 including threaded rod and nut with 2"-12 thread.  These items are under consideration for A+ FC Tube Support Stand, particularly for use in weldment D2000445. Some observations:

  • Thread area seems very small, nuts seem very large - wonder how strong the interface is.
  • Pitch seems cumbersome - minutes may be necessary to install a nut in the middle of the thread.
  • External thread appears to have some small damaged areas. Thread of nut can run over these areas, but with increase in friction.
  • No apparent wobble or loosness of nut due to fit.
  • Nut appears to have a oily coating applied to the outer surfaces, but not to internal threads.

Also ordered were two McMaster offerings which could be used for rust inhibiting conversion coatings (formulations appear to be based on phosphoric acid - more could be learned by investigating specific products). Potential workflow for experimenting:

  1. Learn about the operating instructions for the specific product
  2. Apply the coatings to 1 nut each (preclean needed for outer surfaces?)
  3. Apply the coatings to different regions of the threaded rod

This experiment would allow us to learn about how the conversion coatings may work and behave.

See attached photos and video for more insights.

Attachment 1: IMG_7794.JPG
IMG_7794.JPG
Attachment 2: IMG_7795.JPG
IMG_7795.JPG
Attachment 3: IMG_7796.JPG
IMG_7796.JPG
Attachment 4: IMG_7799.MOV
Attachment 5: IMG_7801.JPG
IMG_7801.JPG
Attachment 6: IMG_7802.MOV
Attachment 7: IMG_7804.JPG
IMG_7804.JPG
  246   Fri Oct 23 12:25:13 2020 Marie K.ProgressBS BRDsBS BRDs for LLO are assembled and tuned

The 5 BS BRDs for installation at LLO are now assembled and tuned in the LLO Optics Lab. The target frequencies are 16.846 Hz and 24.672 Hz (measurements from A. Effler, I will post an alog in the LLO logbook about it).
As per the LHO ones, I tuned them at lower frequencies (-1 or -2%) in order to anticipate the frequency drifts.

For the LHO & LLO bounce modes and for LLO roll modes, I had to put two copper blocks per end at the blade tips in order to reach the target frequency.

There are several factors to take into account:

  1. The measured mass for each type of the copper mass is lighter than what was expected from the design D1500429-v4 (about 0.2 to 0.3 g difference, see table in figure 2)
  2. Somehow I have to use heavier masses than expected from the experiments at CIT (between 0.7 and 1.5 g, see below)
Bounce CIT LLO LHO
Target Freq [Hz] 16.64 16.85 17.79
Mass [g] 9.89 10.43 10.150
k [N/m] 108.11 116.81 126.82
Expected mass* [g] 9.89 9.65 8.65
Diff 0.00 -0.78 -1.50

 

Roll CIT LLO LHO
Target Freq [Hz] 24.34 24.67 26.06
Mass [g] 7.32 7.876 7.247
k [N/m] 171.20 189.27 194.29
Expected mass* [g] 7.32 7.12 6.39
Diff 0.00 -0.75 -0.86

* Extrapolated from the k CIT experimental k.

I would have expected smaller variations in the spring constant.

 

The resulting masses are:

  • H1 Bounce: lightest bounce type (type 8, ~8.6g) + spare mass of 0.595g + washer+screw
  • L1 Bounce: heaviest bounce type (type 14, ~8.8g) + spare mass of 0.595g + washer+screw
  • H1 Roll: heaviest roll type (type 7, ~6.4g) + washer+screw
  • L1 Roll: heaviest roll type (type 7, ~6.4g) + spare mass of 0.55g + washer+screw

Even if the mass kit had the expected weight, I would have not been able to build the BRDs with the masses from the kit (or use several washers of 0.1g).

Attachment 1: LLO_BRDs.jpg
LLO_BRDs.jpg
Attachment 2: Table1__MassKit.png
Table1__MassKit.png
  245   Thu Oct 22 14:45:22 2020 AidanComputingHWS_PADHWS point absorber diagnostic - frame grabber not responding.

The EDTpdv framegrabber is unaccessible following a security update to the HWS computer. The lspci command shows the EDT frame grabber card as a PCI device but the initcam code no longer can connect to it. 

HWS code was backed up to GIT.

Tried updating EDT software to fix this but didn't work. After exhausting options, I'm going to reinstall the OS and try installing the EDT drivers again. Going to try to install Ubuntu 20.04 first (currently running 14.04).

  244   Thu Oct 22 14:33:50 2020 Marie K.ProgressBS BRDsBS BRDs for LHO are assembled and tuned

The 5 BS BRDs for installation at LHO are now assembled and tuned in the LLO Optics Lab. The target frequencies are 17.79 Hz and 26.06 Hz (ref alog 49643).

The masses I had to use for the Bounce mode are slightly heavier than I was expecting (above 10.10g instead of 9.90g). The masses for the Roll mode are in target. I deliberately tuned the BRDs to lower frequencies (between -1% and -2% below the target) to anticipate the drift that shifts the frequencies to higher values. After 2 days, the drifts seem lower than was previoulsy observed at CIT (about 0.1% per day compared to 0.5% per day) with some of the BRDs exhibiting no drift at all. That might be good news? Monitoring will continue over the next weeks.

The Q values fluctuate from measurement to measurement, but they are essentially close to 100.

I am using the LLO B&K vibrometer kit, with an older version of the software that Stuart help me to setup, to measure the resonant frequencies. The air flow from the flow bench is currently off.

Attachment 1: LHO_BS_BRDs.jpg
LHO_BS_BRDs.jpg
Attachment 2: Test_setup.jpg
Test_setup.jpg
Attachment 3: Monitoring_2020_10_22.png
Monitoring_2020_10_22.png
  243   Mon Sep 28 09:31:30 2020 Calum TorrieMechanicsTMDSTMDS FRS TICKET 15312

https://services.ligo-la.caltech.edu/FRS/show_bug.cgi?id=15312

  242   Tue Jun 23 16:18:44 2020 Isabella Dula ProgressBS BRDsFinite Element Analysis on BS BRD blades V5, V6, and simplified rectangular blade

Basis structural analysis was completed on models of the V5, V6, and a simple rectangular version of the BS BRD blades. The modesl were simplified to remove the glue layers. The models include stress analysis for individual layers of the blades, stress analysis for the entire blade, and a linearized stress analysis to examine stress discontinuties within the cross-section of the blade. The stresses analyzed include equivalent, maximum principle, and normal stress. The total deformation was also examined in the models. 

There is also attached an excel summary of the results of the models. 

The initial resuts of these models suggest that there is increased structural weakness in the V6 blade as opposed to the V5 blade. However, the results do not match with theorhetical stress analysis for the blades, so further analysis must be completed to refine the accuracy of the models. 

Attachment 1: structural_analysis_results.xlsx
Attachment 2: blade_rect_basicstructuralanalysis_pass3.wbpj
<?xml version="1.0" encoding="utf-8"?>
<Storage>
  <Project Version="9.1">
    <base-designpoint-name valType="String">0</base-designpoint-name>
    <timestamp-max valType="Int64">13625</timestamp-max>
    <framework-build-version valType="String">19.3.111.0</framework-build-version>
    <external-version-string valType="String">2019 R1</external-version-string>
    <last-saved-utc valType="String">05/18/2020 15:28:54</last-saved-utc>
    <managed-categories valType="String">[report, misc, progress]</managed-categories>
    <logHistory valType="StringArray">
... 478 more lines ...
Attachment 3: blade_rect_basicstructuralanalysis_pass3.wbpj
<?xml version="1.0" encoding="utf-8"?>
<Storage>
  <Project Version="9.1">
    <base-designpoint-name valType="String">0</base-designpoint-name>
    <timestamp-max valType="Int64">13625</timestamp-max>
    <framework-build-version valType="String">19.3.111.0</framework-build-version>
    <external-version-string valType="String">2019 R1</external-version-string>
    <last-saved-utc valType="String">05/18/2020 15:28:54</last-saved-utc>
    <managed-categories valType="String">[report, misc, progress]</managed-categories>
    <logHistory valType="StringArray">
... 478 more lines ...
Attachment 4: bladev6_basicstructuralanalysis_pass3.wbpj
<?xml version="1.0" encoding="utf-8"?>
<Storage>
  <Project Version="9.1">
    <base-designpoint-name valType="String">0</base-designpoint-name>
    <timestamp-max valType="Int64">13951</timestamp-max>
    <framework-build-version valType="String">19.3.111.0</framework-build-version>
    <external-version-string valType="String">2019 R1</external-version-string>
    <last-saved-utc valType="String">05/18/2020 15:41:06</last-saved-utc>
    <managed-categories valType="String">[report, misc, progress]</managed-categories>
    <logHistory valType="StringArray">
... 479 more lines ...
  241   Wed Mar 4 11:32:43 2020 Marie K.ProgressBS BRDsBRD9v5 & BRD10v5 on BS

Right after tuning the BRDv5 9 and 10 (see elog 238), I installed them on the BS and measured the BS transfer function.

In figure 1, there are two consecutive measurements of the roll transfer function. The measurements are very close in time (~1 hour). However, we can see a large difference in the measured Q (lower than a factor 2). It is due to a change of settings in the power spectrum analyzer. First measurement is with auto range OFF and second measurement is with autorange ON. We can see a spike at 24.64 Hz in the blue curve. That corresponds to the moment I changed the settings. Edit: Rich A. just explained to me that the "autorange ON" setting isadjusting the gain as a function of the input voltage it is seeing. THerefore it must be used at the beginning of the experiment to select the gain and chek the noise floor and be turned OFF during measurements.

However, the BRDs are in tune for the roll mode.

Similarly, I measured the bounce mode right after installation (see figure 2) with the auto settings on. The Q of the peak is high compared to the measurements for the other BRDs. Could it be related to the setting change once again? Yes, I have to remeasure the mode.

Attachment 1: BRD9&10_Feb26_roll.png
BRD9&10_Feb26_roll.png
Attachment 2: BRD9&10_Feb26_bounce.png
BRD9&10_Feb26_bounce.png
  240   Fri Feb 28 11:22:06 2020 Marie K.ElectronicsBS BRDsScheme of the electronics for BS measurements

For reference, please find the description of the current installation in the drawing attached.

Attachment 1: BS_electronics_setup_scheme.png
BS_electronics_setup_scheme.png
  239   Thu Feb 27 12:16:22 2020 Marie K.ProgressBS BRDsBRD5_v5 & BRD6_v5 off the BS

Today I removed the BRD5_v5 and BRD6_v5 from the BS suspension, after their 6 months stay.

I measured their tuning right after. The identification of the BRD is not visible (it is written on the blade but hidden in the mount). So I just assigned a new name to the BRDs.

The measurements below confirm that the BRDs are still within range(+/-1%) after 6 months.

=================================

BRD5a_v5 - Bounce Mode

Measurement # Frequency [Hz] Q
4 16.672 118
5 16.703 113
6 16.656 -
Mean 16.68 116

Tuning is 0.1%

BRD5a_v5 - Roll Mode

Measurement # Frequency [Hz] Q
1 24.250 84
2 24.234 86
3 24.234 86
Mean 24.240 85

Tuning is 0.4%

================================

BRD6a_v5 - Bounce Mode

Measurement # Frequency [Hz] Q
1 16.578 53
2 16.531 43
3 16.578 50
Mean 16.562 49

Tuning is 0.8%

BRD6a_v5 - Roll Mode

Measurement # Frequency [Hz] Q
21 24.219 -
22 24.188 67
23 24.141 66
Mean 24.183 67

Tuning is 0.6%

 

  238   Thu Feb 27 11:50:02 2020 Marie K.ProgressBS BRDsBRD9_V5 and BRD10_v5 installed today on BS

BRD9_v5 and BRD10_v5 were remeasured and tuned yesterday (see elog 235).

I double checked the tuning right before installing the BRDs on the BS. Only BRD10 Roll mode was out of the 1% range (F = 24.05 Hz, too low), so I retuned it.

For reference, the final tuning is:

=================================

BRD9_v5 - Bounce Mode

Measurement # Frequency [Hz] Q
57 16.656 78
58 16.656 83
59 16.656 80
Mean 16.656 80

Tuning is 0.2%

BRD9_v5 - Roll Mode

Measurement # Frequency [Hz] Q
60 24.469 59
61 24.453 55
62 24.406 -
Mean 24.443 57

Tuning is 0.4%

================================

BRD10_v5 - Bounce Mode

Measurement # Frequency [Hz] Q
12 16.672 81
13 16.656 69
14 16.688 -
Mean 16.672 75

Tuning is 0.1%

BRD10_v5 - Roll Mode

Measurement # Frequency [Hz] Q
21 24.281 -
22 24.297 -
23 24.250 -
Mean 24.276 -

Tuning is 0.3%

  237   Wed Feb 26 14:40:52 2020 Marie K.SummaryBS BRDsBS Roll mode : BRD5_v5 & BRD6_v5

====== Roll mode

Here is a summary of the BS roll mode survey with BRD5_v5 and BRD6_v6 attached to the suspension for 6 months.
Details of the measurements are attached in the spreadsheet.

The measured transfer functions over time are shown in figure 1. We observed two peaks in the data. This is in agreement with our model if the tuning if the BRD is within 0.1% of the BS roll mode (see T1900846).

  • Trend .  The mean frequency is Hz. The maximum excursion is +% and the minimum is -%.  
  • The Q of the mode is varying from to. Correlation with the frequency variations.

Therefore it seems that the bounce damping is pretty stable over 6 months. We didn't analyze correlations of the variations with environmental factors in the lab.

Measurements after taking off the BRD from BS:

Recall that pre-installation measurements were: 


  236   Wed Feb 26 11:24:35 2020 Marie K.SummaryBS BRDsBS Bounce mode : BRD5_v5 & BRD6_v5

====== Bounce Mode

Here is a summary of the BS bounce mode survey with BRD5_v5 and BRD6_v6 attached to the suspension for 6 months.
Details of the measurements are attached in the spreadsheet.

The measured transfer functions over time are shown in figure 1. We observed a single peak in the data. This is unexpected from our model.

  • There is no trend in the value of the resonant frequency.  The mean frequency is 16.660 Hz. The maximum excursion is +0.4% and the minimum is -0.2%.  
  • The Q of the mode is varying from 60 to 260, with no obvious correlation with the frequency variations. The maximum Q = 260 corresponds to a measurement with a lower amplitude of excitation (see elog xxx).

Therefore it seems that the bounce damping is pertty stable over 6 months. We didn't analyze correlations of the variations with environmental factors in the lab.

Measurements after taking off the BRD from BS:

Recall that pre-installation measurements were:

Attachment 1: BRD5&6_Winter_2019_survey_bounce_last.png
BRD5&6_Winter_2019_survey_bounce_last.png
Attachment 2: Picture1.png
Picture1.png
  235   Wed Feb 26 10:46:53 2020 Isabella Dula ProgressBS BRDsTuning BRD9V5 and BRD10V5

Constructed and tuned the BRD9 and BRD10 V5. The masses on each BRD were:

  Roll Mode mass (g) Bounce Mode mass (g)
BRD9V5 7.31 10.061
BRD10V5 7.309

10.060

 

In the attached excel spreadsheet are the measured frequencies of the BRDs before and after tuning the dampeners. The BRDs were tuned to be with +/- 1% of the ideal value for the dampeners. These frequencies are 16.69 Hz for the bounce mode and 24.34 Hz for the roll mode. Also, the change in frequency over this week-long process was recorded to determine potential drift of the dampeners. 

Attachment 1: BRD9_10_V5.xlsx
  234   Wed Feb 19 13:28:18 2020 StephenItems to BuyGeneralGigE Camera - IR Sensitivity improvements available

Contributors

Stephen, Noah

Summary

It appears that Basler offers GigE cameras with over 2x improved sensitivity to the typical LIGO infrared laser wavelength. This camera is the "acA1300-60gmNIR" and it appears to be 2 to 5 times more sensitive at 1064 nm.

We do not have any of these cameras, and might want to consider getting our hands on one to evaluate its utility in lab settings (or perhaps even in site GigE camera installations).

Detailed Findings

As discussed in the Basler technical documentation, there are 3 different sensors that Basler packages into their cameras.

--> ref. https://www.baslerweb.com/en/vision-campus/camera-technology/nir-cameras/

"The NIR-optimized cameras with NIR-optimized 2 MP (CMV2000) and 4 MP (CMV4000) sensors from CMOSIS, or the 1.3 MP sensor (EV76C661) from e2V, still manage quantum efficiencies close to 40% in the 850nm range. Compared to non-NIR-optimized cameras, this represents a doubling of the sensitivity value at this wavelength."

Examining these sensors closer, the better QE at 1000 nm wavelength is from the 1.3 MP sensor (EV76C661) from e2V. This is notable because LIGO uses 1064 nm wavelength, and the highest QE reported in the datasheets is at 1000 nm.

The camera which uses this sensor is the "acA1300-60gmNIR" which is available in either C-mount lens interface, or CS-mount lens interface.

--> ref. https://www.baslerweb.com/en/products/cameras/area-scan-cameras/ace/aca1300-60gmnir/

Methods

Begin at a given camera's Basler webpage. See example below.

--> ref. https://www.baslerweb.com/en/sales-support/downloads/document-downloads/basler-ace-aca2000-50gmnir-emva-data/

Navigate to "Documents" tab and then to "EMVA Data". There are two attachments in this section:

  • the overview pdf, which gives a rundown of the specs of each camera and the cameras using each sensor
  • the sensor-specific pdf, which gives technical data from the tests collected for a number of articles of each sensor, which corresponds to the indicated sensor

By using the camera-specific pdf, you can identify the quantum efficiency at NIR wavelenghts up to 1000 nm.

Alternatively, if you know you want a specific sensor or specification, you can identify the cameras using the overview pdf, or using the Vision System Configurator

--> ref. https://www.baslerweb.com/en/products/tools/vision-system-configurator/#/selection/camera/

Next Steps

We'll use this log as a starting point to compile the resolutions, sensitivities, and any other parameters for each existing LIGO GigE camera and each possible improved camera.

We will also continue our work in identifying all of the necessary components that could help us construct a mobile setup which enables interchangeability of the various GigE cameras that are at our disposal (WIP).

  233   Wed Feb 19 10:58:46 2020 DulaProgressBS BRDsTuning BRD9V5

Data is in attached excel file.

Attachment 1: BRD9_10_V5.xlsx
  232   Mon Jan 27 14:31:28 2020 Maria GProgressBS BRDsBRD 7 and BRD 8

From January 14 to the 21, Bounce BRD 7's frequency drifted -2.68% which differes from the drift logged on January 14 which was 0.15% on the opposite direction. The Roll of BRD 7's frequency drifted -4.41% which is a big change from the previous logged drift of 0.06%.

The Roll side for BRD 7 on January 21 was much closer to the target frequency. It was logged as 24.36 Hz and the target is 24.34 Hz. The Bounce side was also logged as much closer to the target frequncy. The target is 16.69 Hz and the frequency logged was 16.80 Hz.

The Q factors for both the Roll side and the Bounce side have decreased from the previous entry. Roll decreased from 108 Hz to 100 Hz and Bounce decreased from 87 Hz to 79 Hz.

BRD 8's frequency was tested for the first time. Roll side's frequency was 24.57 Hz. Bounce side's frequency was 17.956 Hz. Roll side is close to the target frequency, but Bounce side is off by 1.26 Hz.

 

The length for the new copper masses will be computed.

Attachment 1: BRD_7_v5_and_BRD_8_v5.xlsx
Attachment 2: Copper_Mass_Lengths.xlsx
  231   Thu Jan 16 16:05:14 2020 Marie K.ProgressBS BRDsExpected Q vs mistuning

To determine the best achievable Q as a function of the mistuning, we studied a simple model with 2 resonant masses (see elog 150). Now we have a model with 3 resonant masses (presented in elog 217, 219, T1900846).

  • Factor 2:

As a reminder, in order to obtain the same transfer function with 2 masses model and the 3 masses model, we need to multiply the mass of the unique BRD by 2 in the model with 2 masses (see figure 1).

Here we assumed the Q of the standalone BRD is 110 and perfect tuning. The resulting Q of the two peaks is around 220. This is in line with the rule of thumb that states "for perfect tuning, the resultant Qs of the bounce and roll modes are approximately 2x the Q of the damper" (Norna's email). However, with the 3 masses model, the resonance of the BS corresponds to the "dip" in the transfer function. At this frequency, the Q value is exactly the Q of the damper. Therefore, there is no factor 2 for the BS mode.

  • Detuning

%% Recalling here some results:

  • mB = 10 g → kB = 110 N/m, 
  • mR = 7.3 g → kR = 171 N/m

Mass ratio is therefore:

  • μB = 10g/27.78 kg = 3.6e-4
  • μR= 7.3g/12.90 kg = 5.7e-4

And we measured:

  • QB ~ 100-110
  • QR ~ 60-80

%%

In T1500271, Brett derives the best achievable Q of a damper for a given mass ratio:

Q = 1/(2*w1/w2*sqrt(3*mu(ii)/(8*(1+mu(ii))^3))); % from equation 12 in T1500271

Comparing these values to our blade v5 measurements for bounce and roll modes, we can see that the actual Qs for the standalone BRDs are a factor 2.5 higher than this ideal case (see figure 2). We therefore updated the study of elog 150 with the new 3 resonant masses model and the Qs of the standalone BRDs equal to 2.5 times the ideal Qs. The results are summarized in graphs 3 and 4. We can see that:

  • The Q of the BS decreases with the increase of the mass ratio. Therefore, the Q of the bounce mode will be higher than the roll mode.
  • The Q of the BS increases with the detuning. 1% mistuning allows to maintain the Qs below 200.
  • The Q of the BRDs decreases with the increase of the mass ratio.
  • The Q of the BRDs decreases with the detuning. At perfect tuning, the Q of the bounce mode is about 225.

So we could aim for a tuning of the BRDs within 1% of the BS resonant frequencies.

Attachment 1: Model_2BRDs_vs_1BRD_v2.png
Model_2BRDs_vs_1BRD_v2.png
Attachment 2: IdealQ.png
IdealQ.png
Attachment 3: Best_Q_vs_mistuning_3masses.png
Best_Q_vs_mistuning_3masses.png
Attachment 4: Best_Q_vs_mistuning_3masses_BRD.png
Best_Q_vs_mistuning_3masses_BRD.png
  230   Tue Jan 14 14:59:47 2020 Maria GProgressBS BRDsTuning BRD 7

Overnight the Bounce mode of BRD7_v5 drifted 0.16% which is surprising if we compare it to the drift from the previous month which was lower and in the other direction. The Roll mode of BRD7_v5 drifted by 0.06% and also in the positive direction. As mentioned in elog 154 we want to tune the BRD to 16.69 Hz for Bounce and 24.34 Hz for Roll.

We tuned the bounce side at 0.25% from the target at 16.731 Hz. We noticed that the Q factor increased from 60 to 90 during the tuning. Could it be related to the tightness of the screw? 

We tuned the Roll side 0.07% from the target at 24.347 Hz. Again we see an increase in the Q factor to over 100.

Attachment 1: BRD_7_v5.xlsx
  229   Tue Jan 14 14:47:33 2020 Maria GProgressBS BRDsBRD 7 and 8 Drift Witnesses

Today we added the masses for the Roll of BRD 8:

4.043+2.351+screw= 7.293g

Quote:

We built 2 BRDs with version 5 of the blade before the break to be used as witnesses for the drift over the next weeks. BRD 8 is made with a baked blade from last summer. We didn't have time to properly tune the BRDs, but they can still be used as witnesses.

BRD 7 BRD 8 (baked)
Bounce: 4.228+4.229+screw= 9.269g Bounce: 4.221+2.42+2.35+screw= 9.843g
Roll: 4.173+2.366+screw= 7.360g Missing screw to complete

Here are the measurment results for BRD 7 before and after the break. We see a negative drift for both modes. For the roll mode the drift is -0.1% and the bounce mode is -0.3%.

 

 

  228   Mon Jan 13 14:36:32 2020 Maria GProgressBS BRDsBRD 7 and 8 Drift Witnesses

We built 2 BRDs with version 5 of the blade before the break to be used as witnesses for the drift over the next weeks. BRD 8 v5 is made with a baked blade from last summer. We didn't have time to properly tune the BRDs, but they can still be used as witnesses.

BRD 7 v5 BRD 8 v5 (baked)
Bounce: 4.228+4.229+screw= 9.269g Bounce: 4.221+2.42+2.35+screw= 9.843g
Roll: 4.173+2.366+screw= 7.360g Missing screw to complete

Here are the measurment results for BRD 7 v5 before and after the break. We see a negative drift for both modes. For the roll mode the drift is -0.1% and the bounce mode is -0.3%.

 

Attachment 1: BRD_7_v5.xlsx
  227   Fri Jan 10 11:53:22 2020 Craig WorleyProgressModal TestingRelayed measurements of VMD

CIT Modal lab is currently experimenting using a mirror relay to measure the VMD at a 45 degree angle to the mounting plate. The rational behind this is the desire to measure the approximate resonance frequency of both the first and second mode in one experiment. Future work will include validating this measurement with tuning of the VMD to make sure we are correctly identifying the modes.

Attachment 1: laseralignment.jpg
laseralignment.jpg
  226   Fri Dec 13 13:06:36 2019 Jordan ProgressClean and BakeSmall ABO Controller Upgrade

Below are the temperature v. time plots from the new Platinum series contoller/software. The profile ran was a 2.5 hour ramp to 94°C, 2.5 hour soak, and 2.5 hour ramp down to room temp. The controller had a maximum overshoot of ~4°C. The Extech datalogger was also used to track temperature, with a maximum overshoot of ~ 5°C.

Attachment 1: Controller.png
Controller.png
Attachment 2: Datalogger.png
Datalogger.png
Attachment 3: Profile_2-test4.xlsx
  225   Fri Dec 13 10:57:23 2019 Marie K.ProgressBS BRDsSurvey of BRD_v5 drifts on BS (Fall 2019)

Attached are the measurements of the bounce roll transfer function of the dummy BS with BRD5 and BRD6 (v5 blades) fastened to the pum.

BRD5_v5 was built on July 16 (elog 165) and BRD6_v5 was built on August 21 (elog 198) after baking (elog 204). It was found that BRD5_v5 stopped drifting at the end of August (elog 204). They were attached on the suspension on August 28th after tuning (elog 202). The date of the measurements is recalled in the table attached, with the latests ones from December 4th, 2019. There are some measurements missing from early November unfortunately (issues with the spectrum analyzer saving data/USB key problems).

Nevertheless, we can clearly see that the BS resonances are still damped over the time period of 3 months, meaning that BRDs are still well in tune with the BS resonances. It means that they stopped drifting or that the drift is now negligeable.

Note: the measurements 206, 207, 208, 212 were taken with a lower amplitude of excitation (250 mV instead of 1V) to check if a high amplitude excitation was not introducing non-linear effects. It might be that the coherence is lower for these measurements and there are therefore less reliable.

 

Attachment 1: BRD5&6_Fall_2019_survey_bounce.png
BRD5&6_Fall_2019_survey_bounce.png
Attachment 2: BRD5&6_Fall_2019_survey_roll.png
BRD5&6_Fall_2019_survey_roll.png
Attachment 3: August2019_monitoring.xlsx
  224   Thu Dec 12 11:16:03 2019 StephenProgressGeneralSR3 ROC Actuator photos

StephenA, various others :)

Sharing a link to photos of SR3 ROC Actuator efforts in various labs at CIT - others involved in various measurements are welcome to add their own photos here.

https://photos.app.goo.gl/1BvwYRREtNP1CCpP7

Related documents:

Notes about the lab setups:

  • In the lab at CIT, we make use of a type K thermocouple (Accuglass p/n 100770 with Hot Junction flat washer lug termination, aka 1/4-20 washer). Accuglass non-controlled print is attached.
    • This auxiliary thermocouple is not used in situ. This caused confusion as documented in FRS Ticket 10063. There is an internal temperature sensor native to the heater which is used.

 

Attachment 1: image3.JPG
image3.JPG
Attachment 2: accuglass_type_k_thermocouple_with_flat_washer_6-103011.pdf
accuglass_type_k_thermocouple_with_flat_washer_6-103011.pdf
  221   Thu Dec 5 11:19:11 2019 Isabella DulaProgressBS BRDsBRD4 V4 Measurements

 

See attached file for measurements. 

Attachment 1: BRD4_v4_DEC19.xlsx
  220   Thu Nov 21 10:53:59 2019 StephenLab InfrastructureClean and BakeElectrical Check of new Epoxy Curing "Wait Test" Oven

RichA, StephenA, LizN 12 November 2019 (catching up after-the-fact)

A new simple oven has been procured which will be dedicated to dirty operations like the "Wait Test" (in other words, elevated temperature checks that proportions and procedures of epoxy mixing have been adequate mixing of epoxy, such as EP30-2 - ref. T1300322 section I.5).

This oven has been property tagged and Liz was planning to help input into the PCS system - C32329 is the property tag but currently there is a conflict in PCS; here is the conflicting entry https://services.ligo-la.caltech.edu/Inventory/history.php?recid=1590&invtype=cit_noncap

Rich has reviewed the electrical characteristics of the toaster (Power = 1150 W per manufacturer spec, so current drawn will be < 10 A, and therefore no issue plugging into standard wall outlet) and has checked dis/continuity of ground and power at the plug. Rich has given the OK to proceed using this oven.

Photos show oven, property tag, and receipt.

 

Attachment 1: IMG_5994.JPG
IMG_5994.JPG
Attachment 2: IMG_5995.JPG
IMG_5995.JPG
Attachment 3: IMG_6008.JPG
IMG_6008.JPG
  219   Tue Nov 19 09:36:56 2019 Marie K.ProgressBS BRDs3 masses model check

Following up elog 217, I want to make sure that the 2 BRD model is correct. Here is a comparison of the model with the 2 identical BRDs compared to the simple model with 1 BRD (the mass is twice the mass of an individual BRD but the Q is unchanged). In both cases the BRDs are tuned at the exact resonant frequency of the BS mode.

We can see that the models are perfectly overlapping. Hence, the 2 BRD model is in agreement with the 1 BRD model (see figure 1)

For the Roll mode, the poles and zeros of the 2 BRD model are:

  • z = {-104662,   -0.1031 +24.3398i,  -0.1031 -24.3398i,  -0.1031 +24.3398i,  -0.1031 -24.3398i}
  • p = {  -0.0547 +24.6018i,  -0.0547 -24.6018i,  -0.1031 +24.3398i,  -0.1031 -24.3398i,  -0.0513 +24.0809i,  -0.0513 -24.0809i}
  • We can see that 1 pair of poles and zeros is cancelling out.

The poles and zeros of the 1 BRD model are:

  • z = {-0.1031 +24.3398i,   -0.1031 -24.3398i}
  • p = {  -0.0547 +24.6018i,  -0.0547 -24.6018i,  -0.0513 +24.0809i,  -0.0513 -24.0809i}

**

If the frequencies of the BRDs are symetrically detuned from the BS resonant frequency in the 2 BRD model, the poles and zeros are:

For 1 % detuning:

  • p = { -0.1042 +24.5832i,   -0.1042 -24.5832i,   -0.1021 +24.0964i,   -0.1021 -24.0964i}
  • z = { -0.0782 +24.6978i,   -0.0782 -24.6978i,   -0.0564 +24.3377i,   -0.0564 -24.3377i,   -0.0746 +23.9869i,   -0.0746 -23.9869i}

For 0.1% detuning :

  • p = { -0.1032 +24.3641i,   -0.1032 -24.3641i,   -0.1030 +24.3154i,   -0.1030 -24.3154i}
  • z = { -0.0551 +24.6028i,   -0.0551 -24.6028i,   -0.1023 +24.3397i,   -0.1023 -24.3397i,   -0.0517 +24.0799i,   -0.0517 -24.0799i}

The resulting transfer function is presented in figure 2.

Attachment 1: Model_check.png
Model_check.png
Attachment 2: Model_check_detuning.png
Model_check_detuning.png
  218   Fri Nov 15 10:45:18 2019 Craig WorleyProgress Data Normalization Script and VMD Testing Document @ CIT

Updated T1900341 Matlab code with additional comments to allow for more ease in changing the code for individual needs. Also updated VMD document T1800474 to clarify the conclusions in mass shift, and fix some grammatical errors. Continued modal testing of Mirror relay in hopes of determining feasibility of using this mounting set up.

  217   Fri Nov 8 13:47:22 2019 Marie K.GeneralBS BRDsModel with 2 BRDs

We want to compare the model transfer function with the measurements on BS.

We built a model with two BRDs attached to the main mass. Model is provided in the attachment. Details of the computation will be given in a note later. We implemented the model in Matlab. 

Back in July (alog 177), we tuned the BRD frequencies right before mounting them on BS. We installed BRD1 and BRD3 (alog 181) with the following properties (alog 177, alog 164, alog 141):

Mode BRD1_v4 Bounce BRD1_v4 Roll BRD3_v4 Bounce BRD3_v4 Roll
Mass (g) 4.936 3.063 4.865 3.207
f (Hz) 16.693 24.26 16.66 24.32
Q 100.4 117.9 154.4 137.4
Accuracy 0.016% -0.34% -0.18% -0.097%

Plugging these properties in our new 2 BRDs model, we obtain the transfer function in Figure 1 and Figure 2. For comparison, the 1 BRD model with only BRD1(3) is shown on the same graph. The tables below show a summary of the model outputs.

Bounce F1 F2 F3 Q1 Q2 Q3
Model 2 BRDs 16.530 16.678 16.837 247 122 239
Model BRD1 16.588 16.795   202 191  
Model BRD3 16.567 16.783   267 337  
Measurement (#34) 16.55 16.64 16.75 108   172

 

Roll F1 F2 F3 Q1 Q2 Q3
Model 2 BRDs 24.051 24.289 24.583 231 128 274
Model BRD1 24.116 24.485   193 283  
Model BRD3 24.144 24.518   257 275  
Measurement (#34) 24.034   24.507 126   160

The 2 BRD model generates three poles but in the present case we observe only "2 peaks", as we have seen in the measurements (see here measurements #33 and #34). We need some larger detuning to see "3 peaks". See for example figure 3 with respectively  -0.1Hz detuning for BRD1 and +0.1Hz detuning for BRD3. The Q are ~ a factor 2 lower than expected from the model.

Attachment 1: Model3BRD_bounce_elog217.png
Model3BRD_bounce_elog217.png
Attachment 2: Model3BRD_roll_elog217.png
Model3BRD_roll_elog217.png
Attachment 3: Model3BRD_roll_detuned_elog217.png
Model3BRD_roll_detuned_elog217.png
Attachment 4: BRD_3masses_elog217.m
%% Use the results from the Model_with_2parallel_BRDs.nb
% based on the model from Norna T1600259
% MK November 2019

clc; clear all; close all;
 
% BS and damper parameters
% primary oscillator with damping
 
% BS Bounce mode
... 136 more lines ...
  216   Thu Nov 7 11:52:01 2019 Luis SanchezElectronicsSatellite Box - OSEMTest Procedure and Data Sheet

Luis

I just added a document to DCC the describe the electrical characteristics from the Aplus Satellite Amplififer E1900355. Also a test procedure was preparer to check all connections on the Aplus Satellite Amplifier E1900245.

Attachment 1: E1900355.pdf
E1900355.pdf E1900355.pdf E1900355.pdf
Attachment 2: E1900245-v2.pdf
E1900245-v2.pdf E1900245-v2.pdf E1900245-v2.pdf E1900245-v2.pdf E1900245-v2.pdf E1900245-v2.pdf E1900245-v2.pdf
  215   Fri Nov 1 10:40:11 2019 Craig WorleyProgress Matlab Script for Comparing Excited and Non-excited Data

Finalized revisions for version two of the Matlab script responsible for intaking B&K files and self-normalizing the different tests to allow for direct comparisons. This program's purpose is to allow for comparisons of different tests in which the excitation level is varied, or binning needs to be completed before comparisons can be made. Outputs include the normalized plot, a plot illustrating the effect of binning selection, and then a file output showing details pertaining 3dB and resonance frequency. More details can be found at T1900341, along with a demonstration video to the posted in the next week.

  214   Thu Oct 31 16:54:45 2019 Marie K.SummaryBS BRDsSummary of the BRDs results with version 4 of the blades

BRDs with the version 4 of the blades have been tested on stand-alone version and in the BS suspension during the summer. Here is a summary of our findings with the references to the corresponding elogs.

Stand alone:

Drifts measured on two BRDs before baking over 20 and 40 days (elog 186). The variation of the resonance frequencies are reported in the table below:

 

BRD1

BRD2

BRD4

 

Bounce

Roll

Bounce

Roll

Bounce

Roll

Time lapse dt [days]

19

41

5

Drift df [Hz]

0.145

0.417

0.218

0.184

0.160

0.196

Drift df [%]

0.87

1.71

0.76

1.31

0.94

0.81

Drift Rate [mHz/day]

7.63

21.96

5.32

4.49

32.00

39.20

Drift Rate [%/day]

0.09

0.04

0.03

0.02

0.18

0.16

  •  The BRD2 modes seem to stabilize after a month of steady increase (see figure 1). The rate of frequency drift is about few hundreds of ppm a day.  The drift rate of BRD1 is of the same order but the measurements were stopped before it stabilized.
  • After baking (elog 197), the BRD4 is measured over 5 days. The rate of the frequency drift dramatically increases (by a factor 4). It might be partly due to the fact that the BRD was mounted after being backed and there is a relaxation in the mount.
  • The value of the Qs is around 150 (elog 177)
  • Reference for all the measurements can be found at: T1900569

On the dummy BS suspension:

1 - Resonance frequencies

The two BRDs were installed on the dummy BS for a month. Unfortunately we are missing some of the measurements because some data got corrupted (see spreadsheet attached). Therefore the analysis is only performed over 16 days (8 days) for the bounce (roll) mode.

For each mode (bounce and roll), two peaks are observed around the resonance where we expected to resolve three peaks. The three peaks would be the main BS resonance as well as one peak per BRD. We might need to increase the scan resolution (see figure 2 and 3).
The frequency of the peaks do not match the frequency of the BRDs measured alone. The frequency of the minima in between the resonance peaks is close to the resonance frequency of the BS for respectively the bounce and the roll modes. The shift in the BS resonance frequency due to the added mass is negligible.

We observed a steady drift of the resonance frequencies over time. The frequencies are increasing by few hundreds of ppm per day, see below and figures 4 and 5.

Summary of BS Bounce frequencies

Time lapse dt [days]

16.00

16.00

Initial mistuning [%]

0.84

-0.34

Final mistuning [%]

0.64

-0.61

Drift df [Hz]

0.034

0.044

Drift df [%]

0.21

0.26

Drift Rate [mHz/day]

2.125

2.75

Drift Rate [%/day]

0.01

0.02

We are observing a change of 0.2% in the resonance frequency over the 16 days. This is above the requirements that we set at 0.1% tuning. We didn’t observe a stabilization in the drift.

Summary of BS Roll frequencies

Time lapse dt [days]

8.00

8.00

Initial mistuning [%]

1.26

-0.69

Final mistuning [%]

1.12

-0.87

Drift df [Hz]

0.034

0.045

Drift df [%]

0.14

0.18

Drift Rate [mHz/day]

4.25

5.625

Drift Rate [%/day]

0.02

0.02

We are observing a change of 0.1% in the resonance frequency over the 8 days. We didn’t observe a stabilization in the drift, so this is likely to exceed our requirements.

When remeasured stand-alone after being uninstalled from the BS suspension, we established that the BRDs frequencies drifted of about 0.5% after being for a month on the suspension (elog 203).

2 - Q factor

The resonance amplitudes, corresponding to the quality factor of the modes, fluctuate over time without a distinguishable pattern. However, it seems that for each mode the two resonances vary together, in particular for the roll modes.

The mean of the bounce mode Q is 140 (168) for the 16.66 Hz (16.69 Hz) resonance. The mean of the bounce mode Q is 115 (152) for the 24 Hz (24.5 Hz) resonance. Q are lower than expected according to the model (we expected Q~200).

Conclusion:

We observed a steady drift of the resonance frequencies of the BRDs over time, when stand alone or on the dummy BS. The frequencies are increasing by few hundreds of ppm per day. We see a stabilization in the drifts after about a month in the lab. The drift is slightly lower when the BRDs are mounted on the BS suspension compared to the stand alone BRDs in the lab. This could confirm that the excitation measurements cause some of the drift and we need to revise the method. We have no evidence that the baking process reduces the frequency drifts.

The measurements of the quality factor shows that the peaks may not be quite resolved, and that we are underestimating the Qs of the modes on the suspension. However, the value of the Qs in the stand-alone measurements is already promising.

 

Attachment 1: BRD1_drifts.png
BRD1_drifts.png
Attachment 2: Bounce_August2019_final.png
Bounce_August2019_final.png
Attachment 3: Roll_August2019_final.png
Roll_August2019_final.png
Attachment 4: BS_bounce_august_drift.png
BS_bounce_august_drift.png
Attachment 5: BS_roll_august_drift.png
BS_roll_august_drift.png
Attachment 6: August2019_monitoring.xlsx
  213   Fri Oct 18 09:42:03 2019 Luis SanchezGeneralA-plusAir Puge valve for HAM7 L1, H1 attemp number 2.

Luis, Rich, Don:

After getting some feedback from Chandra, we agree to revise the documents D1900114 and D1900116. This revision is to allocate the air purge valve on the other side of the chamber, now changing the valve from D6 Flange to D2 Flange. The documents are being modified by Don. As seen in the image, we are also adding the "HDS" on the suspension lines for better identification. The seismic db25 connectors are being allocated to D3 Flange and D3-F11 and D3-F112. The same modifications are being made to document D1900116.

Attachment 1: After_getting_some_feedback_from_Chandra.pdf
After_getting_some_feedback_from_Chandra.pdf
  212   Thu Oct 10 13:01:21 2019 Luis SanchezPlanGeneralAir Puge valve for HAM7 L1, H1

Luis:

A modification on documents D1900114 and D1900116 is on the works, and the changes will be modified by Don, as soon as he has the time. This new change is required because we need a one port for the Air Purge Valve at HAM7. This valve will help to bring the Chamber to room conditions -not under pressure-. Squeezer Tip-Tiilts that are located on D6 Flange (IC1, IC2) will change their location to D8 Flange, which is located at the top of the chamber. D8 will be a  flange with 12 dsubs with 25 pins each.

Attachment 1: Air_Purge_Valve_location_HAM7.pdf
Air_Purge_Valve_location_HAM7.pdf
  211   Wed Oct 9 16:20:01 2019 Luis SanchezElectronicsSatellite Box - OSEMInitial test from D1900217 pcb Suspension Satellite Amplifier

Luis:

Just started to test the pcb D1900217-v2, the voltage noise level appears normal and in range, see following charts. I need to take the Instrument level noise and add this to the charts.

 

Attachment 1: Noise_Test_from_LED_measured_at_R48.pdf
Noise_Test_from_LED_measured_at_R48.pdf Noise_Test_from_LED_measured_at_R48.pdf Noise_Test_from_LED_measured_at_R48.pdf
  210   Fri Oct 4 15:15:22 2019 Luis SanchezElectronicsA-plusSuspension Satellite Amplifier Chassis

Luis:

I assembled two A-Plus Suspension Satellite chassis, and also I seriliazed the pcb boards D1900217 for these D1900089 units, the numbers are S1900540 and S1900541. In the near future I will add all information on the e-traveler, for now I need to do electrical test on these units.

Attachment 1: A-plus_Suspension_Satellite_Chassis_1900089.pdf
A-plus_Suspension_Satellite_Chassis_1900089.pdf
  209   Wed Oct 2 09:53:31 2019 Luis SanchezElectronicsVariosPUM Driver Quote, Binary Output, Squeezer Wiring, and HDS

Luis:

I tried to generate a Sunstone quote for PUM Driver D070483 pcb board, but for some reason the web site was not generating the quote numbers. I asked Patrick from Sunstone and he helped us to generate the quote, now the ODB++ and gerber files are located in Sunstone server for future reference, quote number SQW-52105.

I serialized the pcb boards for Binary Output chassis, and I completed the electrical assembly of 5 units. I still need to serialized the chassis and add all information to the e-traveler. Another step that these chassis require is to do electrical test, also I need to find or to order some 1U handles, screws, etc.

I modified the Squeezer Wiring diagram by removng the VOPO and ZM Suspension elements and I updated this. I will add the removed sections to the A-plus HAM7 chamber.

Created HDS clasification table, We need to define which units will require dither and where these will be located.

Attachment 1: Sunstone_quote_for_the_PUM_Driver_D070483_and_more_.pdf
Sunstone_quote_for_the_PUM_Driver_D070483_and_more_.pdf Sunstone_quote_for_the_PUM_Driver_D070483_and_more_.pdf Sunstone_quote_for_the_PUM_Driver_D070483_and_more_.pdf
  208   Wed Sep 25 13:28:29 2019 Luis SanchezElectronicsVariosChassis, Test, etc

I started to assemble chassis front panel and rear panel for Suspension Satellite Amplifier unit, while pcb board is been manufactured. Suspension Satellite Amplifier rear panel was modified per Rich's comments, I need to center Local Diagnostics outputs.

I continue assembling Binary Output chassis, I found some front and rear panels and also some assembled pcbs that we will use for A+. I still need to connect the led and power on some chassis. total of 5.

Bram contacted us, because he needs PUM driver, this unit is missing a pcb document only schematics, odb and gerber files were found on DCC . I asked Sustone and Patrick mentioned to me that they can manufacture the pcb with ODB files. I updated schematic D070483 on DCC to include the driver board. 

A test was performed on Top Coil Driver with pcb D0902747, appears that circutry is susceptible to oscillations. Could not find the issue. We are thinking that the behaviour described by Dave Hoyland on received email, might be due to a load capacitance due to a long cable connection on Coil outputs, and also to op amp configuration, we saw that design use a front end unit that has high bandwidth 10MHz and unit at the back end has only 1MHz bandwidth. Rich suggested that we could use an OP-06 (slower device) in the front end (600khz bandwidth) in place of the front end op-amp.

 

Attachment 1: Preparing_received_panels_for_chassis_assembly.pdf
Preparing_received_panels_for_chassis_assembly.pdf Preparing_received_panels_for_chassis_assembly.pdf Preparing_received_panels_for_chassis_assembly.pdf Preparing_received_panels_for_chassis_assembly.pdf
  207   Tue Sep 17 11:02:59 2019 Luis SanchezElectronicsA-plusChassis List and getting materials for Binary Input and Output chassis

Started to collect some items for Binary Input and Binary Output chassis, these can be see in below image. Also created document that list the preliminary chassis count for the UK electronics. This list needs to be update since the Tip-Tilts have evolve to a HDS.

Attachment 1: Created_a_chassis_list_for_the_UK_Suspension_Electronics_E1900273.pdf
Created_a_chassis_list_for_the_UK_Suspension_Electronics_E1900273.pdf Created_a_chassis_list_for_the_UK_Suspension_Electronics_E1900273.pdf
  206   Wed Sep 11 11:25:13 2019 Luis SanchezElectronicsA-plusAssembly Drawing A+ Suspension Satellite Amplifier

I just created an assembly drawing for the A+ Suspension Satellite Amplifier on 1U chassis.

Attachment 1: assembly_drawing_for_Sat-Amp.pdf
assembly_drawing_for_Sat-Amp.pdf
  205   Mon Sep 9 13:49:00 2019 Luis SanchezElectronicsA-plusTriple Coil Driver panels, assembly drawing and cable drawings.

After some digging on Jay Heefner old computer files, I finally found the panels for the Triple Coil drivers, I added some panel numbers and I save documents into DCC, the numbers are D1900419, and D1900418.

I also created an assembly drawing since D1001242 only shows a picture of chassis without any description, see below image.

I created a cable drawing for the cables need to connect the monitor board and to suspension satellite amplifier.

Attachment 1: D1001242_eLog.pdf
D1001242_eLog.pdf D1001242_eLog.pdf
  204   Wed Aug 28 11:20:05 2019 Andy R.ProgressBS BRDsBRD Net Drifts

Total drifts of the BRDs. BRD4 and BRD6 were baked. Baking appears to reduce v4 blade drift but increases v5 blade drift. Additionally, BRD5 has definitely stopped drifting and has settled at a max drift of around ~0.6%. 

  BRD4 Roll BRD4 Bounce BRD5 Roll BRD5 Bounce BRD6 Roll BRD6 Bounce
f Drift (Hz) 0.196 0.16 0.14 0.13 0.29 0.36
Percent Drift 0.82% 0.94% 0.56% 0.78% 1.2% 2.2%

 

Attachment 1: BRD4_Roll.png
BRD4_Roll.png
Attachment 2: BRD4_Bounce.png
BRD4_Bounce.png
Attachment 3: BRD5_Roll.png
BRD5_Roll.png
Attachment 4: BRD5_Bounce.png
BRD5_Bounce.png
Attachment 5: BRD6_Roll.png
BRD6_Roll.png
Attachment 6: BRD6_Bounce.png
BRD6_Bounce.png
  203   Wed Aug 28 11:04:00 2019 Andy R.ProgressBS BRDsBRD1 and 3 Updates

BRD1 and BRD3 Resonance peaks after being in BS Suspension for ~1 month. Note that the BRD1 roll mode sagged a lot and the peaks were did not look sharp, so drift is probably unreliable. 

  BRD1 Roll BRD1 Bounce BRD3 Roll BRD3 Bounce
Measured f(Hz) 24.39 16.59 24.45 16.73
Drift in f (Hz) from 7/31 0.133 -0.10 0.075 0.133
Percent Drift 0.55% -0.62% 0.45% 0.55%

There appears to be less total drift when left in the BS suspension when compared to the BRDs that were measured more frequently. This seems to confirm the suspicion that excitation measurements cause some of the drift. 

  202   Wed Aug 28 10:52:06 2019 Andy R.ProgressBS BRDsBRD Tuning

BRD5 and BRD6 are mounted in the BS suspension for testing and were tuned to the following resonant frequencies. Note BRD5 Roll mass was changed to: 4.23g+2.27g+long screw = 7.357g.

  BRD5 Roll BRD5 Bounce BRD6 Roll BRD6 Bounce
Measured Frequency (Hz) 24.412 16.704 24.256 16.656
Difference from Resonance (Hz) 0.072 0.014 -0.084 -0.034
Percent Difference 0.3% 0.08% -0.35% -0.2%
  201   Tue Aug 27 14:59:16 2019 Luis SanchezElectronicsA-plusSuspension Satellite Amplifier pcb test procedure

Luis:

I took some data from D1900217 and created a test procedure. Simulation results are almost in agreement with test data. Limited by SR785 when taking noise spectrum in low frequency specially between 1-10Hz.

Attachment 1: E1900245-v1_082719.pdf
E1900245-v1_082719.pdf E1900245-v1_082719.pdf E1900245-v1_082719.pdf E1900245-v1_082719.pdf E1900245-v1_082719.pdf
ELOG V3.1.3-